LIF SPECTROSCOPY OF ASYMMETRIC DEUTERATED METHOXY

ILIA J. KALINOVSKI^{a}, \underline{X. L. CHEN}, WAI Y. FAN, C. BRADLEY MOORE, Chemical Sciences Division of the Lawrence Berkeley National Laboratory and Department of Chemistry of UC Berkeley, California 94720 and Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

Laser-induced fluorescence spectroscopy of the Ã^{2}A_{1} ↔ X̃^{2}E transition of asymmetrically deuterated methoxy radicals, CHD_{2}O and CH_{2}DO, has been studied. From the origin of the Ã^{2}A_{1} state to 4100 cm^{−1} above it, 75 fluorescence excitation bands that involve modes ν_{2}, ν_{3}, ν′_{5}, ν′′_{5}, ν′_{6}, ν′′_{6} and their combinations have been observed and assigned for CHD_{2}O. Harmonic frequencies, diagonal and off-diagonal anharmonicities have been determined for most of the modes of Ã^{2}A′ CHD_{2}O. A smaller set of bands has been recorded and vibrations assigned for CH_{2}DO. Ten vibrational bands of CHD_{2}O and six of the CH_{2}DO are rotationally analyzed. An effective rotational Hamiltonian is proposed for the ground state, which combines the spin-orbit interaction with a special effective term that is attributable to the removal of the X̃^{2}E vibronic degeneracy induced by the aymmetric deuteration. The LIF results provide a selection of intermediate states to choose from for our ongoing Stimulated Emission Pumping and fluorescence depletion spectroscopy work that is a prelude to studying the state-selective unimolecular reaction dynamics of the electronic ground state of methoxy.

aNovellus Systems, Inc., 2730 Junction Ave., San Jose, CA 95134